# Learning Representations for Independence Testing

### v2

@@ E.4 Independence Testing with MMD
where we have defined a $U$-statistic $U_{1}=\frac{1}{m(m-1)}\sum_{i\neq j}h((x_{i},y_{i}),(x_{j},y_{j}))$ and a remainder term $R_{1}=\frac{1}{m}\sum_{i}h((x_{i},y_{i}),(x_{i},y_{i}))$. For $U_{1}$, we have immediately that $\operatorname*{\mathbb{E}}U_{1}=\operatorname*{\mathbb{E}}h((X,Y),(X^{\prime},Y^{\prime}))=:\mu_{1}$. Noting that both $R_{1}$ and $\mu_{1}$ are necessarily in $[0,\nu^{2}]$, the overall error from this term is therefore
@@
so that $\operatorname*{\mathbb{E}}\left(1-N_{3}^{(\sigma)}/m^{2}\right)=\frac{3}{m}-\frac{2}{m^{2}}$. Moreover, if $\sigma$ and $\sigma^{\prime}$ almost agree except that $\sigma^{\prime}_{k}=\sigma_{l}$ and $\sigma^{\prime}_{l}=\sigma_{k}$, then $\lvert N_{3}^{(\sigma)}-N_{3}^{(\sigma^{\prime})}\rvert\leq 2m$: the only $(i,j)$ which are potentially affected are those of the form $(\cdot,k)$ or $(\cdot,l)$. Using a version of McDiarmid’s inequality for uniform distributions of permutations (Lemma E.1) then gives us that with probability at least $1-\delta_{3}^{\prime}$ over the choice of $\sigma$,
@@
A single transposition changes changes $N_{2}^{(\sigma)}$ by no more than $4m$, and so by Lemma E.1 we obtain that with probability at least $1-\delta_{2}^{\prime\prime}$,
@@
Lemma E.1 (McDiarmid’s inequality for uniform permutations).
@@
Proposition E.2.
Suppose we have samples =$\mathbb{Z}=\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ drawn iid from $\mathbb{P}_{XY}$, and let ${\boldsymbol{\sigma}}=\{\sigma_{1},...,\sigma_{p}\}$ be a set of $p$ permutations sampled uniformly from $\mathcal{S}_{m}$, the permutation group of $[m]$. We define the action of ${\boldsymbol{\sigma}}$ on samples $\mathbb{Z}$ as ${\boldsymbol{\sigma}}\mathbb{Z}=\{(X_{i},Y_{\sigma_{\ell}(i)})\}_{\begin{subarray}{c}i\in[m],\ell\in[p]\end{subarray}}$. Then for the biased HSIC and MMD estimators we have
@@
$\operatorname*{\mathbb{E}}\left[\widehat{\operatorname{MMD}^{2}}(\mathbb{Z},{\boldsymbol{\sigma}}\mathbb{Z})\mid\mathbb{Z}\right]=\widehat{\operatorname{HSIC}}(\mathbb{Z}).$ (22)
@@
$\operatorname*{Var}[\widehat{\operatorname{MMD}^{2}}(\mathbb{Z},{\boldsymbol{\sigma}}\mathbb{Z})]=\operatorname*{Var}[\operatorname*{\mathbb{E}}[\widehat{\operatorname{MMD}^{2}}(\mathbb{Z},{\boldsymbol{\sigma}}\mathbb{Z})\mid\mathbb{Z}]]+\operatorname*{\mathbb{E}}[\operatorname*{Var}[\widehat{\operatorname{MMD}^{2}}(\mathbb{Z},{\boldsymbol{\sigma}}\mathbb{Z})\mid\mathbb{Z}]]\geq\operatorname*{Var}[\widehat{\operatorname{HSIC}}(\mathbb{Z})].$
@@
Figure 12 shows that the higher variance of the permuted MMD estimator negatively impacts its overall test power compared to HSIC. Also, including more permutations seems to decrease the overall variance, although never lower than that of HSIC.
In practice, the kernels learned by MMD-D and HSIC-D may not correspond, and so Proposition E.2 is inapplicable. That said, we observe the same phenomenon in experiments. Figure 13 shows estimates of the asymptotic variance of MMD and HSIC along a training trajectory. For permuted MMD we consider both a single shuffling of the data (MMD-full), as well as a split shuffling (MMD-split) where we use half the data for our joint distribution sample, and the other half to permute for our product-of-marginals sample. We note that the initial variance of MMD-split is substantially higher than that of MMD-full, which is much higher than the variance of HSIC. MMD-split/full also exhibit greater final variances, particularly at larger batch sizes.
@@ E.5 SNR Pitfall
Recent work by Ren et al. (2024) argues a so-called pitfall of the HSIC signal-to-noise ratio paradigm. They identify a corner case whereby, when the bandwidth of kernel $k$ or $l$ approaches 0, ignoring the threshold term causes the SNR objective $J_{w/o}=\widehat{\operatorname{HSIC}}_{b}/\hat{\sigma}_{\mathfrak{H}_{1}}$ to differ from the true asymptotic power objective $J_{w/}=\left(\widehat{\operatorname{HSIC}}_{b}-r/m\right)/\hat{\sigma}_{\mathfrak{H}_{1}}$ by a factor of $-(m-1)$, where $r$ denotes the asymptotic threshold. We argue that this is not a cause for concern, and in some cases, ignoring this is even preferred.
@@
First, the behavior of the SNR objective in the bandwidth limit tells us nothing about the actual global maximum. Their argument tells us that $J_{w/o}=-(m-1)J_{w/}$ when the bandwidth $w_{X}\to 0$ and for a specific estimate of $r$ [^12] They use $r=\operatorname*{\mathbb{E}}[m\widehat{\operatorname{HSIC}}_{b}]$; however, this does not imply that $J_{w/o}$ explodes as $w_{X}\to 0$ since the SNR objective is optimized at a fixed sample size $m$. In fact, we expect $J_{w/}$ to be negative in this limit, since arbitrarily small bandwidths $w_{X}=0^{+}$ almost certainly produce poor test power, which arises from a negative asymptotic power expression (6). Second, their argument seems entirely dependent on the choice of estimators. For instance, if we use $\widehat{\operatorname{HSIC}}_{u}$ rather than $\widehat{\operatorname{HSIC}}_{b}$, and take the asymptotic threshold to be $\operatorname*{\mathbb{E}}[m\widehat{\operatorname{HSIC}}_{u}]=m\operatorname{HSIC}(\mathbb{P}_{x}\times\mathbb{P}_{y})=0$, we get that $J_{w/o}=J_{w/}$.
@@
As further evidence, we plot both $J_{w/}$ and $J_{w/o}$ at varying bandwidths $w_{X}$ on the ISA dataset used in Ren et al. (2024). We use the same settings as they do: $m=250$, $d=3$, $\theta=\pi/10$, $w_{Y}=1.0$. Results are shown in Figure 14. For very small bandwidths, $J_{w/o}$ does not explode and is not the global maximum. Moreover, the actual maximum agrees relatively well between $J_{w/}$ and $J_{w/o}$ at larger sample sizes $m$. We were unable to reproduce their Figure 1.
@@
In fact, Figure 14 (a) shows that ignoring this threshold term may actually be preferred. Consider the limiting behavior as the bandwidth $\omega_{X}$ goes to infinity. In this regime, the gram and centered gram matrices are respectively $K|_{\omega=\infty}=\mathbf{1}\mathbf{1}^{T}$ and $K_{c}|_{\omega=\infty}=\mathbf{O}$, which tells us that $\widehat{\operatorname{HSIC}}_{b}=\text{Tr}[K_{c}L]/m=0$. Since $m\widehat{\operatorname{HSIC}}_{b}$ is degenerate here, its asymptotic variance also goes to 0 as does its $1-\alpha$ quantile. Thus, both $J_{w/}$ and $J_{w/o}$ go to 0 in this bandwidth limit. Now consider a difficult problem where the power at initialization is roughly less than 0.5, meaning $J_{w/}$ is approximately less than 0. Then the maximum of $J_{w/}$ is erroneously this bandwidth limit.
@@
Figure 14: Plots of $J_{w/o}$ (blue) and $J_{w/}$ (orange) at very small bandwidths $\omega_{X}$ with sample size $m$. For the threshold estimate, we use the .95.95.95.95-quantile of the Gamma approximation. The objective maximum is indicated by stars.
E.6 SNR with Threshold
Although maximizing the HSIC signal-to-noise objective $J_{\operatorname{HSIC}}$ is sufficient for learning powerful tests, it may also be beneficial to estimate the threshold-dependent term in the asymptotic power expression of Equation 6. Adopting the same notation as before, we define the SNR with threshold expression as
@@ E.6 SNR with Threshold
where $\Psi$ is the problem-dependent cdf described in Proposition A.2. Estimating $\tilde{J}_{\operatorname{HSIC}}$ proves troublesome, largely due to $\Psi^{-1}$ depending on both the variables $X,Y$ as well as kernels $k,l$. Additionally, it needs to be differentiated with respect to those kernel parameters. We opt to approximate the null distribution $\Psi$ with a Gamma distribution $F_{\nu,\theta}$ with shape and scale parameters
@@
We compare the performance of HSIC and NDS tests maximized with and without the threshold-dependent term in Figure 15. For HSIC, including this term seems to hurt the overall test power for harder problems. This is expected, since the null distribution is difficult to accurately estimate. We use two approximations here: one being the Gamma approximation, and two being the series expansion used to estimate the derivative of the Gamma cdf with respect to the shape parameter. On the other hand, NDS with the threshold term yields higher test power. Again, this is not too surprising as, unlike for HSIC, the null distribution here follows the CLT and thus is easily approximated given enough samples.
@@
Figure 15: Power vs test size based on NDS and HSIC signal-to-noise objectives, both with (blue) and without (red) estimating the threshold-dependent term. Our proposed methods NDS (12) and HSIC-D (13) correspond to nds-w/ and hsic-w/o, respectively. We consider both HDGM-8 and HDGM-10, and use the same training configuration as specified in Section 6.
E.7 Why do Permutation Tests deviate from Asymptotics?
As Figure 6 suggests, the asymptotics (i.e. tests based on the asymptotic null and alternative distributions) may not explain permutation tests well. Figure 16 illustrates why this might be: we observe strong dependency between the sample statistic $\hat{T}(\mathbb{X},\mathbb{Y})$ and its rejection threshold, estimated via permutation.
@@ E.7 Why do Permutation Tests deviate from Asymptotics?
Figure 16: Plot of the NDS statistic $\hat{T}$ evaluated on a sample $(\mathbb{X},\mathbb{Y})$ from RatInABox, versus its rejection threshold estimated with 200 permutations. Each point corresponds to a separate test set $(\mathbb{X},\mathbb{Y})$ of size $m=1000$, but is evaluated with the same critic function $f$. The results in (a) use a critic that maximizes the NDS test power, as in Section 5. The critic in (b) directly maximizes InfoNCE. The red line is $y=x$, which is the rejection boundary; points above this line (colored yellow) are tests that reject the null hypothesis, while points below (purple) do not.

### v3

@@ F.4 Independence Testing with MMD
where we have defined a $U$-statistic $U_{1}=\frac{1}{m(m-1)}\sum_{i\neq j}h((x_{i},y_{i}),(x_{j},y_{j}))$ and a remainder term $R_{1}=\frac{1}{m}\sum_{i}h((x_{i},y_{i}),(x_{i},y_{i}))$. For $U_{1}$, we have immediately that $\operatorname*{\mathbb{E}}U_{1}=\operatorname*{\mathbb{E}}h((X,Y),(X^{\prime},Y^{\prime}))=:\mu_{1}$. Noting that both $R_{1}$ and $\mu_{1}$ are necessarily in $0,\nu^{2}]$, the overall error from this term is therefore
@@
so that $\operatorname*{\mathbb{E}}\left(1-N_{3}^{(\sigma)}/m^{2}\right)=\frac{3}{m}-\frac{2}{m^{2}}$. Moreover, if $\sigma$ and $\sigma^{\prime}$ almost agree except that $\sigma^{\prime}_{k}=\sigma_{l}$ and $\sigma^{\prime}_{l}=\sigma_{k}$, then $\lvert N_{3}^{(\sigma)}-N_{3}^{(\sigma^{\prime})}\rvert\leq 2m$: the only $(i,j)$ which are potentially affected are those of the form $(\cdot,k)$ or $(\cdot,l)$. Using a version of McDiarmid’s inequality for uniform distributions of permutations (Lemma˜F.1) then gives us that with probability at least $1-\delta_{3}^{\prime}$ over the choice of $\sigma$,
@@
A single transposition changes changes $N_{2}^{(\sigma)}$ by no more than $4m$, and so by Lemma˜F.1 we obtain that with probability at least $1-\delta_{2}^{\prime\prime}$,
@@
Lemma F.1 (McDiarmid’s inequality for uniform permutations).
@@
Proposition F.2.
Suppose we have samples $\mathbb{Z}=\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ drawn iid from $\mathbb{P}_{XY}$, and let ${\boldsymbol{\sigma}}=\{\sigma_{1},...,\sigma_{p}\}$ be a set of $p$ permutations sampled uniformly from $\mathcal{S}_{m}$, the permutation group of $[m]$. We define the action of $\boldsymbol{\sigma}$} on samples $\mathbb{Z}$ as ${\boldsymbol{\sigma}}\mathbb{Z}=\{(X_{i},Y_{\sigma_{\ell}(i)})\}_{\begin{subarray}{c}i\in[m],\ell\in[p]\end{subarray}}$. Then for the biased HSIC and MMD estimators we have
@@
[…]
@@
[…]
@@
Figure˜12 shows that the higher variance of the permuted MMD estimator negatively impacts its overall test power compared to HSIC. Also, including more permutations seems to decrease the overall variance, although never lower than that of HSIC.
In practice, the kernels learned by MMD-D and HSIC-D may not correspond, and so Proposition˜F.2 is inapplicable. That said, we observe the same phenomenon in experiments. Figure˜13 shows estimates of the asymptotic variance of MMD and HSIC along a training trajectory. For permuted MMD we consider both a single shuffling of the data (MMD-full), as well as a split shuffling (MMD-split) where we use half the data for our joint distribution sample, and the other half to permute for our product-of-marginals sample. We note that the initial variance of MMD-split is substantially higher than that of MMD-full, which is much higher than the variance of HSIC. MMD-split/full also exhibit greater final variances, particularly at larger batch sizes.
@@ F.5 SNR Pitfall
Recent work by Ren et al. (2024) argues a so-called pitfall of the HSIC signal-to-noise ratio paradigm. They identify a corner case whereby, when the bandwidth of one of kernel $k$ or $l$ approaches 0, ignoring the threshold term causes the SNR objective $J_{w/o}=\widehat{\operatorname{HSIC}}_{b}/\hat{\sigma}_{\mathfrak{H}_{1}}$ to differ from the true asymptotic power objective $J_{w/}=\left(\widehat{\operatorname{HSIC}}_{b}-r/m\right)/\hat{\sigma}_{\mathfrak{H}_{1}}$ by a factor of (m1)-(m-1), where $r$ denotes the asymptotic threshold. We argue that this is not a cause for concern, and in some cases, ignoring this is even preferred.
@@
First, the behavior of the SNR objective in the bandwidth limit tells us nothing about the actual global maximum. Their argument tells us that $J_{w/o}=-(m-1)J_{w/}$ when the bandwidth $w_{X}\to 0$ and for a specific estimator $r=\operatorname*{\mathbb{E}}[m\widehat{\operatorname{HSIC}}_{b}]$. This does not imply, however, that $J_{w/o}$ explodes as $w_{X}\to 0$ since the SNR objective is optimized at a fixed sample size $m$, and even in the bandwidth limit $J_{w/o}$ can not be -$\infty$. The latter observation is true because $\widehat{\operatorname{HSIC}}_{b}$ is bounded and our variance estimates $\hat{\sigma}^{2}_{\mathfrak{H}_{1}}$ are also bounded away from zero. Therefore even though $J_{w/o}$ is high when $w_{X}\searrow 0$, this value is not necessarily the global maximum. Second, their argument seems entirely dependent on the choice of estimators. For instance, if we use $\widehat{\operatorname{HSIC}}_{u}$ rather than $\widehat{\operatorname{HSIC}}_{b}$, and take the asymptotic threshold to be $\operatorname*{\mathbb{E}}[m\widehat{\operatorname{HSIC}}_{u}]=m\operatorname{HSIC}(\mathbb{P}_{x}\times\mathbb{P}_{y})=0$, we get that $J_{w/o}=J_{w/}$ in this regime.
@@
As further evidence, we plot both $J_{w/}$ and $J_{w/o}$ at varying bandwidths $w_{X}$ on the ISA dataset used in Ren et al. (2024). We use the same settings as they do: $m=250$, $d=3$, $\theta=\pi/10$, $w_{Y}=1.0$. Results are shown in Figure˜14. For very small bandwidths, $J_{w/o}$ does not explode and is not the global maximum. Moreover, the actual maximum agrees relatively well between $J_{w/}$ and $J_{w/o}$ at larger sample sizes $m$. We were unable to reproduce their Figure 1.
@@
Additionally, Figure˜14 (a) shows that ignoring this threshold term may actually be preferred. Consider the limiting behavior as the bandwidth $\omega_{X}$ goes to infinity. In this regime, the Gram and centered Gram matrices are respectively $K|_{\omega=\infty}=\mathbf{1}\mathbf{1}^{T}$ and $K_{c}|_{\omega=\infty}=\mathbf{0}$, which tells us that $\widehat{\operatorname{HSIC}}_{b}=\text{Tr}[K_{c}L]/m=0$. Since $m\widehat{\operatorname{HSIC}}_{b}$ is degenerate its variance and $1-\alpha$ quantile are zero, and so both $J_{w/}$ and $J_{w/o}$ are also 0 in this bandwidth limit. [^13] Recall that our variance estimate in $J$ adds a small positive constant for stability, preventing $J$ from being undefined. Now consider a difficult problem where the power at initialization is less than 0.5, meaning $J_{w/}$ is less than 0. Then the maximum of $J_{w/}$ is erroneously this bandwidth limit.
@@
Figure 14: Plots of $J_{w/o}$ (blue) and $J_{w/}$ (orange) at very small bandwidths $\omega_{X}$ with sample size $m$. For the threshold estimate, we use the .95.95-quantile of the Gamma approximation. The objective maximum is indicated by stars.
F.6 SNR with Threshold
Although maximizing the HSIC signal-to-noise objective $J_{\operatorname{HSIC}}$ is sufficient for learning powerful tests, it may also be beneficial to estimate the threshold-dependent term in the asymptotic power expression of Equation˜6. Adopting the same notation as before, we define the SNR with threshold expression as
@@ F.6 SNR with Threshold
where $\Psi$ is the problem-dependent cdf described in Proposition˜A.2. Estimating $\tilde{J}_{\operatorname{HSIC}}$ proves troublesome, largely due to $\Psi^{-1}$ depending on both the variables $X,Y$ as well as kernels $k,l$. Additionally, it needs to be differentiated with respect to those kernel parameters. We opt to approximate the null distribution $\Psi$ with a Gamma distribution $F_{\nu,\theta}$ with shape and scale parameters
@@
We compare the performance of HSIC and NDS tests maximized with and without the threshold-dependent term in Figure˜15. For HSIC, including this term seems to hurt the overall test power for harder problems. This is expected, since the null distribution is difficult to accurately estimate. We use two approximations here: one being the Gamma approximation, and two being the series expansion used to estimate the derivative of the Gamma cdf with respect to the shape parameter. On the other hand, NDS with the threshold term yields higher test power. Again, this is not too surprising as, unlike for HSIC, the null distribution here follows the CLT and thus is easily approximated given enough samples.
@@
Figure 15: Power vs test size based on NDS and HSIC signal-to-noise objectives, both with (blue) and without (red) estimating the threshold-dependent term. Our proposed methods NDS (12) and HSIC-D (13) correspond to nds-w/ and hsic-w/o, respectively. We consider both HDGM-8 and HDGM-10, and use the same training configuration as specified in Section˜6.
F.7 Why do Permutation Tests deviate from Asymptotics?
As Figure˜6 suggests, the asymptotics (i.e. tests based on the asymptotic null and alternative distributions) may not explain permutation tests well. Figure˜16 illustrates why this might be: we observe strong dependency between the sample statistic $\hat{T}(\mathbb{X},\mathbb{Y})$ and its rejection threshold, estimated via permutation.
@@ F.7 Why do Permutation Tests deviate from Asymptotics?
Figure 16: Plot of the NDS statistic $\hat{T}$ evaluated on a sample $(\mathbb{X},\mathbb{Y})$ from RatInABox, versus its rejection threshold estimated with 200 permutations. Each point corresponds to a separate test set $(\mathbb{X},\mathbb{Y})$ of size $m=1000$, but is evaluated with the same critic function $f$. The results in (a) use a critic that maximizes the NDS test power, as in Section˜5. The critic in (b) directly maximizes InfoNCE. The red line is $y=x$, which is the rejection boundary; points above this line (colored yellow) are tests that reject the null hypothesis, while points below (purple) do not.